OpenStudy (anonymous):

Maya had $27. She spent all the money on buying 3 burgers for $x each and 2 sandwiches for $y each. If Maya had bought 2 burgers and 1 sandwich, she would have been left with $11. A student concluded that the price of each burger is $5 and the price of each sandwich is $6. Which statement best justifies whether the student's conclusion is correct or incorrect? Choices: The student's conclusion is correct because the solution to the system of equations 3x + 2y = 11 and 2x + y = 16 is (5, 6). The student's conclusion is incorrect because the solution to the system of equations 3x –

OpenStudy (anonymous):

Maya had $27. She spent all the money on buying 3 burgers for $x each and 2 sandwiches for $y each. If Maya had bought 2 burgers and 1 sandwich, she would have been left with $11. A student concluded that the price of each burger is $5 and the price of each sandwich is $6. Which statement best justifies whether the student's conclusion is correct or incorrect? choices: The student's conclusion is correct because the solution to the system of equations 3x + 2y = 11 and 2x + y = 16 is (5, 6). The student's conclusion is incorrect because the solution to the system of equations 3x – 2y = 11 and 2x – y = 16 is (5, 6). The student's conclusion is correct because the solution to the system of equations 3x + 2y = 27 and 2x + y = 16 is (5, 6). The student's conclusion is incorrect because the solution to the system of equations 2x + 3y = 27 and x + 2y = 16 is (5, 6).

OpenStudy (anonymous):

well is this true 3*5-2*6=11

OpenStudy (anonymous):

15-12=11 not true so go to b

OpenStudy (anonymous):

The answer is B?

OpenStudy (anonymous):

no we have to work it out please give me a little time and i will give you the answer

OpenStudy (anonymous):

Ohhh okay

OpenStudy (anonymous):

it is c

OpenStudy (anonymous):

just tag me when you need help

OpenStudy (anonymous):

alright?

OpenStudy (anonymous):

okay thank you :)